Bob walks 3.5 mph to the store. If he arrives at the store in 17 minutes, how far away is the store?
Found 2 solutions by drk, richwmiller:
Answer by drk(1908)

This is a R * T = D problem; rate x times = distance. The idea in any of these problems is to make sure your units are all the same. They give you a rate in miles per "hour" and a time in "minutes".
We want to turn the minutes into hours; in general turn the smaller unit into the larger unit. Minutes / 60 gives you hours. So, we have
3.5 m/h * 17/60 hr = miles.
We simply multiply 3.5 x 17 and divide by 60 to get .9916666 miles.

Answer by richwmiller(17219)

you have to convert minutes to hours
so 17/60=hours
3.5*17/60